OpenStudy (anonymous):

Matt Johnson takes out a mortgage for $240,000. There is a loan of 30 years at $1,200 per month. This gives a total interest of $182,000. What is the APR using the formula?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

182+240 = 240(1+.%)^30 422/240 = (1+.%)^30 30rt(211/120) = 1 + .% 30rt(211/120) - 1 = .% =abt .01899 APR might be abt 1.9%